OpenStudy (saifoo.khan):

The sum of "z" consecutive integers is 90. Which of the following cannot be the value of "z". 3, 4, 5, 6, 9.

mathslover (mathslover):

it seems that the answer given by saurav is very correct

OpenStudy (saifoo.khan):

i still don't get it. lol.

OpenStudy (anonymous):

maybe first put it like this: \[n+(n+1)+(n+2)+...+(n+(z-1))=90\] and find it from here

OpenStudy (saifoo.khan):

How? Why? @myko

OpenStudy (anonymous):

guys lets work it step by step \[n+1+n+2+...+n+z=90\]\[nz+\frac{z(z+1)}{2}=90\]\[z^2+(2n+1)z=180\]\[z(z+2n+1)=180\] we can find all integers \(z\) that satisfies the last equality...but here we have options and we can plug them into equation and see what happens...

OpenStudy (anonymous):

the expretion would be equal to: n*z+ sum of first (z-1) integers = 90

OpenStudy (anonymous):

@mukushla your sum is of z+1 consecutive integers

OpenStudy (anonymous):

no its z n+1 n+2 ... n+z

OpenStudy (anonymous):

should be nz+z(z-1)/2=90

OpenStudy (saifoo.khan):

So..?

OpenStudy (anonymous):

2nz+z^2-z=180 z^2+(2n-1)z-180=0 and I think just check for possible natural solutions of this

OpenStudy (anonymous):

now u have \[n=\frac{90}{z}-\frac{z+1}{2}\] now plug the values of \(z\) and if the value of \(n\) becomes an integer then that \(z\) is your answer

OpenStudy (saifoo.khan):

@mukushla : Your's and @myko 's methods are same right?

OpenStudy (anonymous):

yes they are

OpenStudy (anonymous):

for example \(z=3\) gives \[n=\frac{90}{3}-\frac{3+1}{2}=30-2=28\]

OpenStudy (anonymous):

yes they are same

OpenStudy (anonymous):

only difference is that i start fomr n and @mukushla from n+1

OpenStudy (saifoo.khan):

So answer is 3?

OpenStudy (anonymous):

it asks: Which of the following cannot be the value of "z". and 3 aparently , is.

OpenStudy (saifoo.khan):

But it's incorrect. :S

OpenStudy (anonymous):

let me complete the solution \[z=3 ---------> n=\frac{90}{3}-\frac{3+1}{2}=28\] \[z=4 ---------> n=\frac{90}{4}-\frac{4+1}{2}=20\] \[z=5 ---------> n=\frac{90}{5}-\frac{5+1}{2}=15\] \[z=6 ---------> n=\frac{90}{6}-\frac{6+1}{2}=11.5\] \[z=9 ---------> n=\frac{90}{9}-\frac{9+1}{2}=5\] so the answer is only z=6

OpenStudy (anonymous):

4and 6 would be the answer

OpenStudy (anonymous):

i think

OpenStudy (saifoo.khan):

Myko's thinking is going the correct way!

OpenStudy (anonymous):

i got it by mine first formula using n like a start integer.

OpenStudy (anonymous):

but aparently there should be no difference in starting from n+1 or n+2 or n+whatever

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

for \(z=4\) \[21+22+23+24=90\]

OpenStudy (rsadhvika):

4 90/4 - 4+1/2 90/4 - 10/4 80/4 is an integer

OpenStudy (saifoo.khan):

4 is incorrect too!

OpenStudy (anonymous):

its impossible ... see my last reply

OpenStudy (saifoo.khan):

The answer is 6. :S

OpenStudy (saifoo.khan):

I'm going to type the solution.. just a sec.

OpenStudy (anonymous):

@saifoo.khan thats right answer is 6...

OpenStudy (saifoo.khan):

How? Why? :D

OpenStudy (anonymous):

man... we already solved it...

OpenStudy (saifoo.khan):

But you were getting 11.5 when we inserted 6.

OpenStudy (anonymous):

well that means there is not 6 consecutive integers such that their sum is 90

OpenStudy (saifoo.khan):

(behind the book) Solution: Plug in numbers! This may take some time, but we will eventually find the answer. Three numbers 20+30+31=90,eliminate A, four numbers 21+22+23+24=90,eliminate b,five numbers 16+17+18+19+20=90,elimnate c.nine numbers 6+7+8+9+10+11+12+13+14=90 elimate e.no six consecutive numbers add up to 90

OpenStudy (saifoo.khan):

@mukushla , got it right!

OpenStudy (saifoo.khan):

I totally get it. Thanks a ton! @mukushla & @myko

OpenStudy (anonymous):

yw :)

OpenStudy (anonymous):

yw